A 25 watt bulb emits monochromatic yellow light of wavelength of 0.57um.calculate the rate of emission of quanta per second

Grade:12th pass

1 Answers

Khimraj
5 years ago
Power of bulb, P = 25 Watt = 25 Js–1
energy of photon 
E = hc/\lambda = 34.87*10-20 J
rate of emission of quanta per second = P/E = 7.2*1019 s-1
Hope it clears.